OpenStudy (anonymous):

Find the distance between the two points. Round to the nearest tenth if necessary. (8,8) (12, 11)

OpenStudy (anonymous):

use distance formula

OpenStudy (anonymous):

5

OpenStudy (anonymous):

http://www.purplemath.com/modules/distform.htm

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

5 units

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

4 over 3 up gives a nice 3-4-5 right triangle

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Find the distance bewtween the two points. Round to the nearest tenth if necessary. (-2, -6) (3,9)

OpenStudy (anonymous):

mark this problem as has a good answer so that other people do not waste time coming to help here.

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

5 over, 15 up \[\sqrt{5^2+15^2}\] \[\sqrt{250}=15.81\] rounded

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

want to help me with more problems?

OpenStudy (anonymous):

satellite, I think you made a mistake. your earlier answer of 5 was correct.

OpenStudy (anonymous):

sure, but make sure to post a different question in a different thread.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

because you marked this thread as has a good answer. People will not come here to help.

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

dh i answered a different question. i hope it was correct!

OpenStudy (anonymous):

oh okay.